OpenStudy (calculusxy):

MEDAL! When asked to find the equation (in slope-intercept form) to perpendicular or parallel line with a given slope, when do we change the y-intercept and when do we keep it the same?

TheSmartOne (thesmartone):

For perpindicular if they give you a point along with the slope, you change the y-intercept otherwise you keep it the same. For parallel, I guess you change it all the time :P

TheSmartOne (thesmartone):

perpendicular*

TheSmartOne (thesmartone):

but if you're not sure, it doesn't hurt to look at the answer choices :P

OpenStudy (calculusxy):

okay

OpenStudy (calculusxy):

can you help me with another question but with dividing radicals?

OpenStudy (calculusxy):

\[\frac{ 9 }{ \sqrt{25} }\]

OpenStudy (calculusxy):

@TheSmartOne

TheSmartOne (thesmartone):

\(\sf\Large \frac{9}{\sqrt{25}} = \frac{9}{\sqrt{25}} \times \frac{\sqrt{25}}{\sqrt{25}} = \frac{9\sqrt{25}}{25}\)